If the position vector of A and B are 3


−→a − 7


−→b − 7


−→c and 5


−→a + 4


−→b + 3−→c , nd −→AB


and determine its magnitude and direction cosines.


1
Expert's answer
2022-03-03T17:43:27-0500
"\overrightarrow{AB}=(5-3, 4-(-7), 3-(-7))"

"\overrightarrow{AB}=(2,11,10)"

"|\overrightarrow{AB}|=\sqrt{2^2+11^2+10^2}=15"

Direction cosines are "\bigg(\dfrac{2}{15}, \dfrac{11}{15}, \dfrac{2}{3}\bigg)."